# Mechanisms vs. Outcomes: Probing for Syntax Fails to Explain Performance on Targeted Syntactic Evaluations

### v1

@@ 4 Results & Discussion
Fig 2 plots PTB test $g^{\text{struct}}_{\phi}$ UUAS across model layers. Best layers for GPT-2 (Radford et al., 2019) match with previous results from the literature (Eisape et al., 2022), validating our probe training procedure. In Appendix E, Fig 13 plots PTB test $g^{\text{head}}_{\phi}$ UAS, and we further show that our syntax probes’ abilities to recover encoded parse tree information in BLiMP sentences correlate well across training objective ($g^{\text{struct}}_{\phi}$ vs. $g^{\text{head}}_{\phi}$, Fig 14) and architecture ($g^{\text{struct}}_{\phi}$ vs. $g^{\text{ortho}}_{\phi}$, Fig 15). Furthermore, the comparable range of UUAS/UAS scores between the PTB test set and BLiMP shows the probes generalize from human-written PTB text to template-generated BLiMP.
@@
Figure 2: Penn Treebank test set structural probe UUAS for each layer of a sample of our models. The star icon for a model indicates the layer with the best test set accuracy that is used for BLiMP evaluation. For most models, this occurs in the first half. Our results for GPT-2 117M (small) and GPT-2 1.6B (XL) closely match those of Eisape et al. (2022), supporting the correctness of our implementation. For T5 3B, although most layers yield low UUAS, the best layer (13) still exceeds 0.7.
Figure 3: Control probes consistently erase contextual information in hidden states, as evidenced by near-zero variance of the previously-contextual word contextual hidden states in the projected representation space.
Figure 4: Simple regression plots for $g^{\text{struct}}_{\phi}$ (top row), $g^{\text{ortho}}_{\phi}$ (middle row), and $g^{\text{head}}_{\phi}$ (bottom row). Each panel is annotated with adjusted $R^{2}$ and the $p$-value of $\beta_{1}$. Per-phenomenon results have Holm-Bonferroni correction. The first column of panels shows that at the full dataset granularity, no probe explains the spread in minimal pairs accuracy with any statistical significance. At per-phenomenon granularity, the second column contains the phenomenon with the lowest $p$-value per probe. We additionally highlight subject-verb agreement (third column) and filler-gap (fourth column) as strongly syntactic tasks with critical edges that we identify in Appendix D. Refer to Appendix F for complete simple regression results.
@@
$\beta_{1}$ $p$-value Adj.$R^{2}$ $\beta_{1}$ $p$-value Adj.$R^{2}$ Stat $p$-value
@@
Subject Verb Agr -0.108 1.0 0.007 -0.047 1.0 0.128 5.236 0.243
Filler Gap Dependency -0.028 1.0 -0.033 0.033 1.0 -0.05 0.546 1.0
Table 1: Comparison of simple and multiple regression statistics with $x_{1}=$ $g^{\text{struct}}_{\phi}$ UUAS and $x_{2}=$ $g^{\text{ctrl}}_{\phi}$ $\rho_{s}$. Per-phenomenon $p$-values are Holm-Bonferroni corrected and bold text indicates statistical significance ($p$ < 0.05). While the simple regression model is a better fit for anaphor agreement than the multiple regression model, there is no clear signal across any other phenomena. Refer to Table 5 in Appendix F for complete simple-multiple regression statistics for all phenomena.
@@
$\beta_{1}$ $p$-value Adj.$R^{2}$ $\beta_{1}$ $p$-value Adj.$R^{2}$ Stat $p$-value
@@
Subject Verb Agr -0.119 1.0 0.014 -0.059 1.0 0.132 5.183 0.251
Filler Gap Dependency -0.1 1.0 -0.027 -0.058 1.0 -0.049 0.402 1.0
Table 2: Comparison of simple and multiple regression statistics with $x_{1}=$ $g^{\text{ortho}}_{\phi}$ UUAS and $x_{2}=$ $g^{\text{ctrl}}_{\phi}$ $\rho_{s}$. Per-phenomenon $p$-values are Holm-Bonferroni corrected and bold text indicates statistical significance ($p$ < 0.05). While the simple regression model is a better fit for anaphor agreement than the multiple regression model, there is no clear signal across any other phenomena. The higher capacity of $g^{\text{ortho}}_{\phi}$ relative to $g^{\text{struct}}_{\phi}$ does not translate to improved regression fit, suggesting that its additional probe dimensions are extra. Refer to Table 6 in Appendix F for complete simple-multiple regression statistics for all phenomena.
@@
$\beta_{1}$ $p$-value Adj.$R^{2}$ $\beta_{1}$ $p$-value Adj.$R^{2}$ Stat $p$-value
@@
Subject Verb Agr -0.205 1.0 0.041 -0.221 1.0 0.025 0.56 1.0
Filler Gap Dependency -0.125 1.0 -0.028 -0.15 1.0 -0.055 0.256 1.0
Table 3: Comparison of simple and multiple regression statistics with $x_{1}=$ $g^{\text{head}}_{\phi}$ UAS and $x_{2}=$ $g^{\text{ctrl}}_{\phi}$ $\rho_{s}$. Per-phenomenon $p$-values are Holm-Bonferroni corrected and bold text indicates statistical significance ($p$ < 0.05). The LRT in the first row indicates the superiority of the multiple regression model, but its adjusted $R^{2}$ is still very weak. Refer to Table 7 in Appendix F for complete simple-multiple regression statistics for all phenomena.
No significant fit on full BLiMP.
The first column of Fig 4 shows OLS simple regression result plots at the full BLiMP dataset granularity. Each of the three syntax probes demonstrates a minimal effect size with none reaching statistical significance. $g^{\text{head}}_{\phi}$ has no measurable effect at all. Its UAS values generally exceed the UUAS values of other probes, reflecting successful learning of relation directionality, but there is insufficient stratification along the y-axis for any predictive power.
Table 1 ($g^{\text{struct}}_{\phi}$), Table 2 ($g^{\text{ortho}}_{\phi}$), and Table 3 ($g^{\text{head}}_{\phi}$) compare simple and multiple regression statistics. For $g^{\text{head}}_{\phi}$, the significant $p$-value of the LRT conveys that considering non-syntactic signal from activations provides better fit on full BLiMP than relation direction-oriented signal alone, but the adjusted $R^{2}$ is still very weak.
@@
The second, third, and fourth columns of panels in Fig 4 show simple regression result plots for anaphor agreement, subject-verb agreement, and filler-gap dependencies respectively. Out of the three syntax probes and all phenomena (full regression tables in Appendix F), the only statistically significant correlations are $g^{\text{struct}}_{\phi}$ and $g^{\text{ortho}}_{\phi}$ on anaphor agreement. Probes show poor fit on minimal pairs accuracy for all other phenomena. Strikingly, probes show no predictive power even for phenomena like subject-verb agreement and filler-gap dependencies whose behavior we might expect to be closely reliant on syntactic representations.
However, the anaphor agreement result warrants caution. Fig 5 shows the surprising result that the control probe, which is trained on a non-morphosyntactic task, can be highly correlated to the BLiMP irregular forms morphosyntactic task. Non-syntactic probes can predict syntactic benchmark performance. In the absence of broad systematic correlations, one-off probe/task correlations are not evidence of syntactic representations in hidden states being a primary explanatory variable for downstream behavior, even at the per-phenomenon level: probes designed to be unrelated can be correlated to single tasks. Even with Holm-Bonferroni correction, the anaphor agreement fit could be an artifact of multiple statistical testing. As for the full BLiMP results, Table 1 ($g^{\text{struct}}_{\phi}$), Table 2 ($g^{\text{ortho}}_{\phi}$), and Table 3 ($g^{\text{head}}_{\phi}$) present the simple and multiple regression statistics side-by-side. Aside from anaphor agreement, there is no consistent result of the simpler regression model providing a statistically significant better fit than the regression that includes the control. As a further example, $g^{\text{head}}_{\phi}$ encodes hierarchical dominance of the head word, which we would expect to influence NPI licensing; however, Table 7 shows no such fit.
@@
Figure 5: Control $g^{\text{ctrl}}_{\phi}$ unexpectedly achieves statistical significance for predicting the BLiMP irregular forms phenomena, which is a morphosyntactic task.
Moderate mismatch rate in finer-grained test suite evaluation.
@@
Fig 6 shows that for 3 out of the 4 challenging subject-verb agreement and filler-gap paradigms discussed in Section 3, the match rate between $g^{\text{struct}}_{\phi}$ correctly predicting the critical edge and the model resolving the minimal pair correctly is between 40–60% for the majority of models. The scatter plots show a substantial spread in model performance on both axes, which suggests the difficulty of the tasks. However, high Hamming distances between correct prediction binary outcomes mean that errors do not align as much as expected. We conclude that extracting these critical edges, despite its possible relevance to deciding between the words which are different between the minimal pairs (Appendix D), does not in fact predict evaluation success on the minimal pair. Fig 16 yields the same conclusion from $g^{\text{head}}_{\phi}$ probe results.
@@
Figure 6: Challenging subject-verb agreement and filler-gap phenomena critical edge prediction results for $g^{\text{struct}}_{\phi}$. Hamming distance is computed between the binary outcome of the structural probe predicting the critical edge correctly in $S_{acc}$ and the binary outcome of the model resolving the minimal pair correctly. Successful critical edge prediction does not translate to successful minimal pair resolution for 3 out of 4 test suites. The majority of models show mismatches occurring in approximately 40% to 60% of cases (Hamming distance buckets 0.4 to 0.6). Generally, the harder the test suite, indicated by clustering of points lower on the y-axis in the scatter plots, the higher the Hamming distance.
@@
UUAS score distributions largely overlap for correct vs. incorrect minimal pairs.
@@
While cross-model correlations between probing and downstream performance have proved elusive, a natural follow-up investigation for each model is the sentence-level comparison of if there is a statistically significant difference in the UUAS score distribution between the pool of sentences where the minimal pair is evaluated correctly and the pool where it is evaluated incorrectly. For each BLiMP test suite, we run a two-sample $t$-test for each model on these pools of UUAS scores and apply Holm-Bonferroni correction on the test $p$-values. The results in Fig 7 show that only 8 out of 67 BLiMP test suites have statistical evidence for the mean of UUAS scores for correct minimal pairs being higher than that of incorrect minimal pairs for at least 5 models. Yet again, we conclude that for most models and test suites, UUAS score of the probe at the sentence level (mechanism) does not predict if the model will succeed at the corresponding minimal pair evaluation (outcome).
Figure 7: Sentence-level UUAS score distribution $t$-test results. Out of 67 BLiMP test suites, only the 26 shown here have at least one model that has a statistically significant difference in the UUAS score means. The $t$-test provides statistical evidence that the mean of UUAS scores where the minimal pair is evaluated correctly is greater than the mean where it is evaluated incorrectly.
@@ 5 Related Work
Our work builds on the extensive literature examining the syntactic capabilities of language models through probing (Hewitt and Manning, 2019; Clark et al., 2019; Müller-Eberstein et al., 2022; Limisiewicz and Mareček, 2021; Diego-Simón et al., 2024, among others).
Tucker et al. (2021) demonstrate that counterfactual interventions on contextual embeddings—designed to erase information identified by a probe—can affect downstream behavior in predictable ways. However, they also find that such effects are not robust across models or probes. Similarly, Eisape et al. (2022) perform causal intervention on transformer states during autoregressive generation from GPT-2. Here, we conduct a more comprehensive study to understand the connection between probing and downstream syntactic behavior across model types, scale, and probe categories. Ruzzetti et al. (2023) take a different approach from probing to explore the linguistic properties of monolingual BERT Devlin et al. (2019). While our work is currently limited to English, Ruzzetti et al. (2023) compare layer weight matrices across monolingual models to uncover linguistically-motivated typological similarities. Their work can serve as a blueprint for future multilingual probing evaluations since common phenomena across languages may be predictable from probes trained on monolingual models. Prior work (Bommasani et al., 2022; Belinkov, 2022) cautions against inferring causality between probed representations and downstream tasks Ravichander et al. (2021); Elazar et al. (2021); Belinkov (2022). Belinkov (2022) surveys probing research, emphasizing the importance of careful task and control design for meaningful insights.
@@ 6 Conclusion
Targeted evaluations like BLiMP minimal pairs reveal the extent of model syntactic knowledge. High BLiMP scores observed across the open-weight models studied here suggest effective learning of syntactic phenomena. Our work shows that the mechanism of recovering syntactic knowledge from model activations through linear probing does not at all predict targeted evaluation performance. Despite extensive research on probing, further work is necessary to develop a robust understanding of how syntax is represented within models.

### v2

@@ 4 Results & Discussion
Fig 2 plots PTB test $g^{\text{struct}}_{\phi}$ UUAS across model layers. Best layers for GPT-2 (radford2019language) match with previous results from the literature (eisape2022probingincrementalparsestates), validating our probe training procedure. In Appendix E, Fig 13 plots PTB test $g^{\text{head}}_{\phi}$ UAS, and we further show that our syntax probes’ abilities to recover encoded parse tree information in BLiMP sentences correlate well across training objective ($g^{\text{struct}}_{\phi}$ vs. $g^{\text{head}}_{\phi}$, Fig 14) and architecture ($g^{\text{struct}}_{\phi}$ vs. $g^{\text{ortho}}_{\phi}$, Fig 15). Furthermore, the comparable range of UUAS/UAS scores between the PTB test set and BLiMP shows the probes generalize from human-written PTB text to template-generated BLiMP.
Figure 2: Penn Treebank test set $g^{\text{struct}}_{\phi}$ UUAS for each layer of a sample of our models. The star icon for a model indicates the layer with the best test set accuracy that is used for BLiMP evaluation. For most models, this occurs in the first half. Our results for GPT-2 124M and GPT-2 1.6B closely match those of eisape2022probingincrementalparsestates. For T5 3B, although most layers yield low UUAS, the best layer (13) still exceeds 0.7.
@@
Figure 3: Control probes consistently erase contextual information in hidden states, as evidenced by near-zero variance of word contextual hidden states in the projected representation space. Control probes shown here are trained on structural probe best layers.
Figure 4: Simple regression plots for $g^{\text{struct}}_{\phi}$ (top row), $g^{\text{ortho}}_{\phi}$ (middle row), and $g^{\text{head}}_{\phi}$ (bottom row). Each panel is annotated with adjusted $R^{2}$ and the $p$-value of $\beta_{1}$. Per-phenomenon results have Holm-Bonferroni correction. The first column of panels shows that at the full dataset granularity, no probe explains the spread in minimal pairs accuracy with any statistical significance. At per-phenomenon granularity, the second column contains the phenomenon with the lowest $p$-value per probe. We additionally highlight subject–verb agreement (third column) and filler–gap (fourth column) as strongly syntactic tasks with critical edges that we identify in Appendix D.
@@
[…]
@@
Subject–Verb Agr -0.108 1.0 0.007 -0.047 1.0 0.128 5.236 0.243
Filler–Gap Dependency -0.028 1.0 -0.033 0.033 1.0 -0.05 0.546 1.0
(a) $x_{1}=g_{\phi}^{\text{struct}}$ UUAS.
@@
[…]
@@
Subject–Verb Agr -0.119 1.0 0.014 -0.059 1.0 0.132 5.183 0.251
Filler–Gap Dependency -0.1 1.0 -0.027 -0.058 1.0 -0.049 0.402 1.0
(b) $x_{1}=g_{\phi}^{\text{ortho}}$ UUAS.
@@
[…]
@@
Subject–Verb Agr -0.205 1.0 0.041 -0.221 1.0 0.025 0.56 1.0
Filler–Gap Dependency -0.125 1.0 -0.028 -0.15 1.0 -0.055 0.256 1.0
(c) $x_{1}=$ $g^{\text{head}}_{\phi}$ UAS.
Table 1: Comparison of simple and multiple regression statistics. Each subtable corresponds to a different syntax probe as $x_{1}$, and $x_{2}=$ $g^{\text{ctrl}}_{\phi}$ $\rho_{s}$. Per-phenomenon $p$-values are Holm-Bonferroni corrected and bold text indicates statistical significance ($p$ < 0.05). The datasets shown correspond to those in Fig 4. For both $g^{\text{struct}}_{\phi}$ and $g^{\text{ortho}}_{\phi}$, the simple regression model fits anaphor agreement better than the multiple regression model does, but no consistent signal emerges for other phenomena. The higher capacity of $g^{\text{ortho}}_{\phi}$ relative to $g^{\text{struct}}_{\phi}$ does not translate to improved regression fit. For $g^{\text{head}}_{\phi}$, despite the statistical significance of the LRT at the full dataset granularity, the adjusted $R^{2}$ of the multiple regression model is very weak. Table 3 in Appendix F has regression results for all phenomena.
No significant fit on overall BLiMP.
The first column of Fig 4 shows OLS simple regression result plots at the full BLiMP dataset granularity. Each of the three syntax probes demonstrates a minimal effect size with none reaching statistical significance. $g^{\text{head}}_{\phi}$ has no measurable effect at all. Its UAS values generally exceed the UUAS values of other probes, reflecting successful learning of relation directionality, but lack of y-axis stratification prevents predictive power.
Table 1 compares simple and multiple regression statistics. For $g^{\text{head}}_{\phi}$, the significant $p$-value of the LRT conveys that considering non-syntactic signal from activations provides better fit on full BLiMP than relation direction-oriented signal alone. Even so, the adjusted $R^{2}$ is still negligible, showing that neither regression meaningfully explains BLiMP accuracy. For $g^{\text{struct}}_{\phi}$ and $g^{\text{ortho}}_{\phi}$, adding the control probe offers no meaningful improvement.
@@
Out of the three syntax probes and all phenomena, the only statistically significant correlations are $g^{\text{struct}}_{\phi}$ and $g^{\text{ortho}}_{\phi}$ on anaphor agreement. Strikingly, probes show no predictive power even for phenomena like subject–verb agreement and filler–gap dependencies whose behavior we might expect to be reliant on syntactic representations. Refer to the second, third, and fourth columns of panels in Fig 4 for simple regression plots for these select phenomena.
Table 1 shows that aside from anaphor agreement, there is no consistent result of the simpler regression model providing a statistically significant better fit than the regression that includes the control. As a further example, $g^{\text{head}}_{\phi}$ encodes hierarchical dominance of the head word. We would expect this to influence Negative Polarity Item (NPI) licensing, as NPIs must be in the structural scope of their licensor. However, our results do not support this expectation; Table 2(c) reports a corrected $p$-value of 1.0 and an adjusted $R^{2}$ of zero.
The anaphor agreement result warrants caution. Fig 5 shows the surprising result that the control probe, which is trained on a non-morphosyntactic task, can be highly correlated to the BLiMP irregular forms morphosyntactic task. Non-syntactic probes can predict syntactic benchmark performance. Without broad systematic correlations, isolated syntax probe–task fits are not evidence that the syntactic information in hidden states that is recoverable by these probes is a primary explanatory variable for downstream behavior; even probes with no syntactic basis can one-off align with syntactic tasks. The anaphor agreement finding could be an artifact of minimal pair accuracy saturation.
@@
Figure 5: Control $g^{\text{ctrl}}_{\phi}$ (trained on $g^{\text{struct}}_{\phi}$ best layers) unexpectedly achieves statistical significance for predicting the BLiMP irregular forms phenomenon, which is a morphosyntactic task. Tables 4 and 5 in Appendix F have full control simple regression results.
Critical edge predication does not align with minimal pair evaluation.
@@
Fig 6 shows that for 3 out of the 4 challenging subject–verb agreement and filler–gap paradigms, the match rate between $g^{\text{struct}}_{\phi}$ correctly predicting the critical edge and the model resolving the minimal pair correctly is between 40–60% for the majority of models. The wide spread of minimal pair accuracies along the y-axis of the scatter plots underscores the relative difficulty of these paradigms; by contrast, in Fig 4 plots most accuracies cluster well above 0.7. However, high Hamming distances between edge prediction and minimal pair evaluation mean that errors do not align as much as expected. Extracting critical edges, despite relevance to deciding between the words that are different between the minimal pairs (Appendix D), does not in fact predict minimal pair outcome. Fig 16 in Appendix G yields the same conclusion from $g^{\text{head}}_{\phi}$ results.
@@
Figure 6: Challenging subject–verb agreement and filler–gap phenomena critical edge prediction results for $g^{\text{struct}}_{\phi}$. We compute the Hamming distance between the probe’s binary outcome of predicting the critical edge in $S_{acc}$ correctly and the model’s binary outcome of resolving the minimal pair correctly. Successful critical edge prediction does not translate to successful minimal pair resolution for 3 out of 4 paradigms. The majority of models show mismatches occurring in approximately 40% to 60% of cases (Hamming distance buckets 0.4 to 0.6). Harder paradigms, indicated by lower minimal pair accuracies in the scatter plots, have higher Hamming distance.
@@
UUAS score distributions largely overlap for minimal pair outcomes.
@@
While cross-model correlations between probing and downstream performance have proved elusive, a natural follow-up investigation for each model is the sentence-level comparison of if there is a statistically significant difference in the UUAS score distribution between the pool of sentences where the minimal pair is evaluated correctly and the pool where it is evaluated incorrectly. For each BLiMP paradigm, we run a two-sample one-sided $t$-test for each model on these pools of UUAS scores and apply Holm-Bonferroni correction on the test $p$-values. The results in Fig 7 show that only 8 out of 67 BLiMP paradigms have statistical evidence for the mean of UUAS scores for correctly predicted minimal pairs being higher than that of incorrectly predicted minimal pairs for at least 5 models. Yet again, we conclude that for most models and paradigms, UUAS score of the probe at the sentence level (mechanism) does not predict whether the model will succeed at the corresponding minimal pair evaluation (outcome).
@@ 5 Related Work
Our work builds on the extensive literature examining the syntactic capabilities of language models through probing (hewitt-manning-2019-structural; clark2019bert; muller2022probing; Limisiewicz2021orth; diego2024polar, among others).
Causal Analysis.
Recent probing work aims to make causal claims between probed representations and observed model downstream behavior. Causal analysis intervenes in model representations to understand their effects. tucker2021modified demonstrate that counterfactual interventions on contextual embeddings—designed to erase information identified by a probe—can predictably affect downstream behavior. However, they find that such effects are not robust across models or probes. Similarly, eisape2022probingincrementalparsestates perform causal intervention on transformer states during autoregressive generation from GPT-2. arora-etal-2024-causalgym introduce a causal interpretability method benchmark, CausalGym, to evaluate interventions sourced from minimal pairs on next-token predictions. Rather than evaluating causality, we focus on probing’s statistical predictive power.
Figure 7: Sentence-level UUAS score distribution $t$-test results. Out of 67 BLiMP paradigms, only the 26 shown here have at least one model that has a statistically significant difference in the UUAS score means.
Advancing Probing Research.
probing-promises-shortcomings surveys the promises and shortcomings of probing research. Our paper addresses several core issues raised and advances the discussion with new empirical evidence. First, Belinkov emphasizes the importance of controls for confounders. We introduce a control probe trained to recover non-syntactic information and demonstrate that it can show correlations with syntactic benchmarks. Second, he notes probe complexity affects interpretation, with simpler probes (e.g., linear) preferred to avoid the probe independently learning the probing task. Our work uses three linear probes with different objective functions and metrics for robustness. Finally, Belinkov warns that probing experiments may conflate properties of the model with properties of the datasets used for training or probing, making it difficult to disentangle the two. We use different datasets for training probes (Penn Treebank) and evaluation (BLiMP), which helps ensure our findings are not artifacts of a single dataset.
While some shortcomings of probing are already known, systematic studies like ours are necessary to thoroughly test hypotheses. Notably, our scope spans a broader range of models and linguistic properties than prior work, including ravichander2021probingprobingparadigmdoes, who focus on verb tense, subject number, and object number in Natural Language Inference with pre-transformer architectures, and elazar2021amnesicprobingbehavioralexplanation, who apply amnesic probing to sequence tagging tasks with BERT devlin2019bertpretrainingdeepbidirectional.
Future Multilingual Study.
Just within the last few months, we have seen new work substantially advancing BLiMP for multilingual use cases. The largest cross-lingual development is MultiBLiMP 1.0 jumelet2025multiblimp10massivelymultilingual, which has over 128,000 minimal pairs across 101 languages specifically for subject–verb agreement. Furthermore, başar2025turblimpturkishbenchmarklinguistic contribute TurBLiMP, a Turkish minimal pairs benchmark with 16 linguistic phenomena. We emphasize in our work that a model’s ability to solve minimal pairs correctly is a desired linguistic property and critical evaluation tool, which başar2025turblimpturkishbenchmarklinguistic also echo. Much of our evaluation framework is applicable to these new benchmarks: MultiBLiMP uses Universal Dependencies and TurBLiMP includes BLiMP phenomena. As momentum builds towards targeted syntactic analysis for more and lower-resource languages, we believe that future work on our framework towards multilinguality would help identify language-specific probing and downstream performance artifacts. While probing has little predictive power in English, replicating the study could yield novel insights from languages with for instance, richer subject–verb agreement systems than English or greater word order flexibility as seen in Turkish.
@@ 6 Conclusion
Targeted evaluations like BLiMP are widely used to benchmark model syntactic knowledge. Given the importance of minimal pairs, our key contribution is to show systematically that probing fails to reliably extract the latent syntactic knowledge within models that correlates with minimal pair accuracy. The high BLiMP scores observed across the open-weight models studied suggest effective learning of syntactic phenomena, and probing task performance is likewise strong. However, across models, probing performance does not align with BLiMP performance, even in fine-grained settings where such alignment might reasonably be expected.
Returning to our “mechanisms vs. outcomes” framework, the mechanism of recovering syntactic knowledge from model activations through linear probing does not at all predict targeted outcomes. A common question in probing research is whether the probed representations are actually used to produce model outputs, motivating the line of work on causality. However, our negative finding regarding predictive power preempts this causal question. Extensive research on probing and interpretability more broadly has advanced our knowledge, but a robust understanding of how syntax is represented within models remains to be developed.